### v2

# Nonlinear Magnetically Charged Black Holes with Phantom Global Monopoles: Thermodynamics, Geodesics, Quasinormal Modes, and Grey-Body Factors

Abstract
We study the properties of a nonlinear magnetic-charged black hole in the presence of a phantom global monopole. By incorporating nonlinear electrodynamics (NED) and exotic scalar fields, we derive an exact black hole solution and analyze its geometric structure, causal properties, and thermodynamic behavior. We examine how the presence of a phantom global monopole modifies the black hole’s Hawking temperature, entropy, and stability conditions, revealing significant changes in its phase structure. Additionally, we investigate the geodesic motion of test particles. The quasinormal mode (QNM) spectrum is computed using the WKB approximation and Pöschl-Teller potential method, providing insights into the perturbative stability of the system. Furthermore, we analyze the grey-body factors that characterize radiation emission, highlighting their dependence on black hole parameters. Our findings indicate that the interplay between phantom energy, NED, and global monopoles introduces observable deviations in strong-field astrophysical phenomena. These results offer potential signatures for testing modified gravity theories and contribute to a deeper understanding of black hole physics in exotic field environments.
Keywords: Nonlinear electrodynamics, Phantom global monopole, Black hole thermodynamics, Quasinormal modes, Grey-body factors.
1 Introduction
Black holes in modified gravity theories and exotic matter environments exhibit profound deviations from general relativity (GR), offering unique opportunities to probe high-energy astrophysics and quantum gravity effects [1, 2, 3, 4, 5]. While classical black hole solutions in GR are well established, alternative theories predict significant modifications in strong-field regimes. These deviations can manifest in astrophysical observations such as black hole shadows, gravitational waves, and high-energy emissions, providing crucial insights into physics beyond GR. Additionally, the presence of exotic fields and nonlinear electrodynamics (NED) around black holes alters their thermodynamic stability, causal structure, and radiation properties. Investigating these effects is essential for distinguishing modified gravity scenarios from standard astrophysical processes. For a comprehensive review of these modifications and their implications, see [3] and references therein.
Among the exotic fields that may alter black hole physics, phantom energy is of particular interest due to its ability to violate the null energy condition, leading to counterintuitive consequences [6]. Unlike conventional dark energy models, phantom fields have an equation of state parameter less than $-1$, which can significantly impact black hole thermodynamics and stability. Observational data have not ruled out phantom energy [11], making it an active research topic across various branches of physics. Notably, phantom energy accretion reduces black hole mass, potentially violating cosmic censorship and the generalized second law of thermodynamics [12, 13, 14]. In Reissner-Nordström (RN)-like black holes, this process increases the charge-to-mass ratio beyond unity, modifying standard black hole dynamics [15, 16]. Furthermore, phantom fields affect gravitational lensing properties [17], strong lensing effects, and light propagation [18, 19, 20]. Beyond gravitational effects, phantom fields alter black hole thermodynamics by increasing temperature as mass decreases, influencing the evaporation process via Hawking radiation [21]. However, the stability of phantom-dominated black holes remains uncertain, as negative-energy fluxes can introduce instabilities in the event horizon structure [22, 25, 23, 24, 26]. Further studies in various black hole models, including BTZ black holes in commutative [25] and noncommutative [26] spaces, have highlighted the need for a deeper investigation into their thermodynamic fate and role in an accelerating universe.
Another fundamental extension of GR involves NED, which modifies Maxwell’s theory to account for strong-field effects [30, 29]. Such modifications arise naturally in high-energy regimes, particularly around magnetized black holes or intense electromagnetic fields [31, 32]. Several studies indicate that NED can lead to regular black hole solutions, mitigating singularities while preserving essential thermodynamic properties [35, 34, 33]. The thermodynamics of black holes within NED frameworks have been extensively analyzed, revealing rich phase structures, critical behavior, and modified heat capacities [36, 37, 39, 38]. However, when combined with exotic fields such as phantom energy, NED introduces additional complexities affecting black hole stability, causal structure, and accretion dynamics [41, 40]. Recent research has explored magnetically charged black holes in NED coupled to phantom fields, uncovering novel regular solutions and black bounce geometries [42]. Additionally, investigations into rotating and nonlinearly charged black holes with anisotropic matter fields have demonstrated significant deviations from standard GR solutions [43]. These modifications influence observable phenomena, including black hole shadows, gravitational lensing, and quasinormal modes (QNMs), providing potential avenues for testing deviations from classical GR [44]. Lower-dimensional studies, such as those on BTZ black holes, further highlight the unique thermodynamic and geometric properties introduced by NED [39]. As research continues to explore the interplay between NED and unconventional matter fields, these studies offer valuable insights into quantum gravity corrections, alternative compact objects, and broader modified gravity theories.
An intriguing and less-explored aspect of modified gravity is the role of topological defects, particularly global monopoles, in black hole physics. Global monopoles arise from the spontaneous symmetry breaking of a global $O(3)$ symmetry to $O(2)$ and differ from gauge monopoles due to their long-range gravitational effects [45]. Unlike gauge monopoles, which have localized energy densities, global monopoles introduce a conical deficit angle that modifies spacetime on large scales. Their gravitational influence has been widely studied, including effects on vacuum polarization [46], quantum singularities [47], and general relativistic solutions [49, 48]. When coupled with black holes, global monopoles significantly alter event horizon properties, geodesic motion, and thermodynamics, leading to distinct phenomena such as modified Hawking radiation [50, 52, 51, 53, 54], gravitational lensing effects [49, 55], and deviations in accretion processes [57, 56]. The presence of a global monopole also affects black hole shadows and other observational features in various modified gravity models [58, 59]. Furthermore, the thermodynamic behavior of such black holes has been extensively examined, revealing critical phase transitions [61, 60], modified entropy relations [62, 63], and possible Planck-scale remnants [64]. Studies on charged and rotating black holes with global monopoles suggest substantial deviations from classical solutions, influencing stability and causal structure [52, 53, 61]. Additional research on global monopoles in noncommutative geometries has shown modifications in quantum tunneling radiation and black hole evaporation [51, 52, 59]. These findings highlight the complex interplay between global monopoles, exotic matter fields, and nonlinear electrodynamics, offering potential avenues to test deviations from GR through astrophysical observations.
Despite extensive research on phantom energy, nonlinear electrodynamics, and global monopoles individually, their combined impact on black hole physics remains largely unexplored. This study aims to fill this gap by systematically investigating how their interplay alters black hole thermodynamics, stability, and observational features. This paper is structured as follows: Section 2 presents the derivation of the black hole solution and its geometric properties, while Section 3 analyzes its thermodynamic behavior. Section 4 investigates the timelike geodesics and the motion of test particles in the black hole spacetime. Section 5 examines QNMs using the sixth-order WKB approximation, cross-checked with the Pöschl-Teller potential. Finally, Section 6 analyzes grey-body factors and their influence on radiation emission, before we conclude in Section 7 with a summary and future outlook.
2 Spacetime Geometry and Black Mole Metric
We begin by investigating a static, spherically symmetric black hole embedded within a phantom global monopole. This configuration arises due to the spontaneous symmetry breaking of a triplet of phantom scalar fields associated with a global $O(3)$ symmetry. The presence of the global monopole modifies the action, which takes the form [45]:
$S=\frac{1}{16\pi}\int\sqrt{-g}Rd^{4}x+\int d^{4}x\sqrt{-g}\mathcal{L}^{M},$ (1)
where the matter Lagrangian is given by
$\mathcal{L}^{M}=\frac{\xi}{2}\partial_{\mu}\chi^{a}\partial^{\mu}\chi^{a}-\frac{\lambda}{4}\left(\chi^{a}\chi^{a}-\eta^{2}\right)^{2}.\qquad a=1,2,3.$ (2)
Here, $\chi^{a}=\eta h\left(r\right)\frac{x^{a}}{r}$ represents a scalar field triplet, where $h\left(r\right)$ is a dimensionless function, $\eta$ is the energy scale of symmetry breaking, and $\lambda$ is a coupling constant. When $\xi=1$, the setup corresponds to a conventional global monopole, characterized by a scalar field with positive kinetic energy. However, in the case where $\xi=-1$, the kinetic energy becomes negative, leading to the formation of a phantom global monopole.
To analyze a static, spherically symmetric black hole in the presence of a phantom global monopole, we consider the following line element:
$ds^{2}=-B\left(r\right)dt^{2}+A\left(r\right)dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (3)
which describes the gravitational field around the black hole. In this metric background, the field equations for the triplet scalar field $\chi^{a}$ reduce to a single differential equation governing $h(r)$ [47]:
$\xi\frac{h^{\prime\prime}}{A}+\left[\frac{1}{Ar}+\frac{1}{2B}\left(\frac{B}{A}\right)\right]\xi h^{\prime}-\frac{2\xi h}{r^{2}}-\lambda\eta^{2}h\left(h^{2}-1\right)=0.$ (4)
Following the approach in [45], we assume the asymptotic behavior $h(r)\to 1$. Under this assumption, the energy-momentum tensor associated with the global monopole takes the form:
$T_{t}^{t\left(\mathrm{M}\right)}=T_{r}^{r\left(\mathrm{M}\right)}=\frac{\xi\eta^{2}}{r^{2}},\quad T_{\theta}^{\theta\left(\mathrm{M}\right)}=T_{\phi}^{\phi\left(\mathrm{M}\right)}=0.$ (5)
On the other hand, when coupling Einstein gravity with a nonlinear electromagnetic field in the presence of a phantom global monopole, the field equations governing the system take the form:
$G_{\mu}^{\nu}=2\left(\frac{\partial\mathcal{L}^{charge}}{\partial F}F_{\mu\alpha}F^{\alpha\nu}-\delta_{\mu}^{\nu}\mathcal{L}^{charge}\right)+8\pi T_{\mu}^{\nu\left(\mathrm{M}\right)},$ (6)
$\nabla_{\mu}\left(\frac{\partial\mathcal{L}^{charge}}{\partial F}F^{\nu\mu}\right)=0,$ (7)
where the nonlinear electrodynamics (NED) Lagrangian density is given by:
$\mathcal{L}^{charge}=\frac{3m}{\left|Q\right|^{3}}\frac{\left(2Q^{2}F\right)^{3/2}}{\left[1+\left(2Q^{2}F\right)^{4/3}\right]^{2}}.$ (8)
Here, $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$, while $m$ and $Q$ correspond to the mass and magnetic charge of the system, respectively.
Next, we adopt a spherically symmetric spacetime ansatz of the form:
$ds^{2}=-f\left(r\right)dt^{2}+\frac{1}{f\left(r\right)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (9)
where
$f\left(r\right)=1-\frac{2M\left(r\right)}{r}.$ (10)
To describe the electromagnetic field, we employ the Maxwell field ansatz [48]:
$F_{\mu\nu}=\left(\delta_{\mu}^{\theta}\delta_{\nu}^{\phi}-\delta_{\nu}^{\theta}\delta_{\mu}^{\phi}\right)B\left(r,\theta\right)=\left(\delta_{\mu}^{\theta}\delta_{\nu}^{\phi}-\delta_{\nu}^{\theta}\delta_{\mu}^{\phi}\right)Q\sin\theta,$ (11)
which directly leads to:
$F=\frac{Q^{2}}{2r^{4}}.$ (12)
Using the previously derived equations, the time component of Eq. (6) simplifies to:
$-\frac{2}{r^{2}}\frac{dM\left(r\right)}{dr}=-\frac{6mQ^{3}}{\left(r^{3}+Q^{3}\right)^{2}}+8\pi\frac{\xi\eta^{2}}{r^{2}}.$ (13)
Integrating this expression, we obtain the metric function that describes a black hole with a nonlinear electromagnetic field in the presence of a phantom global monopole:
$f(r)=1-8\pi\xi\eta^{2}-\frac{2mr^{2}}{r^{3}+Q^{3}}.$ (14)
In the special case where $\xi\rightarrow 0$, this solution reduces to a black hole with a nonlinear magnetic charge. On the other hand, setting $Q\rightarrow 0$ retrieves the Schwarzschild black hole solution with a phantom global monopole [49]. To investigate the curvature behavior of this spacetime, we compute the Ricci scalar $\mathbf{R}$ and the Kretschmann scalar $\mathbf{K}$, given by:
$\mathbf{R}=\frac{16\pi\eta^{2}\xi}{r^{2}}+\frac{4m\left(6Q^{6}r^{2}-3Q^{3}r^{5}\right)}{\left(Q^{3}+r^{3}\right)^{3}r^{2}},$ (15)
$\mathbf{K}=\frac{48m^{2}\left(2Q^{12}-2Q^{9}r^{3}+18Q^{6}r^{6}-4Q^{3}r^{9}+r^{12}\right)}{\left(Q^{3}+r^{3}\right)^{6}}+\frac{128\pi\eta^{2}m\xi}{r^{2}\left(Q^{3}+r^{3}\right)}+\frac{256\pi^{2}\eta^{4}\xi^{2}}{r^{4}}.$ (16)
Both the Ricci scalar $\mathbf{R}$ and the Kretschmann scalar $\mathbf{K}$ exhibit divergences at $r\rightarrow 0$, signifying the presence of a curvature singularity at the origin. However, in the asymptotic limit $r\rightarrow\infty$, both scalars vanish, implying that the spacetime flattens at large distances.
To further understand the influence of the phantom global monopole on the black hole structure, we plot the metric function $f(r)$ for different values of the symmetry-breaking scale $\eta$. This allows us to examine how the presence of the monopole affects the event horizon and overall spacetime geometry.
(a) $\eta=0.05$
(b) $\eta=0.1$
(c) $\eta=0.15$
Figure 1: Behavior of the metric function $f(r)$ for different values of the monopole parameter $\eta$, with fixed charge $Q=0.5$ and mass $m=1$.
As observed in Figure 1, increasing the symmetry-breaking parameter $\eta$ leads to a more pronounced deviation of the metric function $f(r)$. Specifically, higher values of $\eta$ result in a deeper gravitational potential, shifting the position of the event horizon outward. This is a direct consequence of the monopole’s additional energy contribution, which effectively modifies the gravitational field.
Moreover, the presence of the parameter $\xi$ (implicitly affecting $\eta$) influences the global behavior of the solution, indicating that a sufficiently strong monopole field can significantly alter black hole thermodynamics and causal structure. In particular, the limiting behavior at large $r$ remains consistent, reinforcing the asymptotic flatness of the spacetime. However, near the black hole horizon, the modifications become more pronounced, potentially impacting observational signatures such as gravitational lensing and black hole shadows.
3 Thermodynamics
In this section, we focus on exploring the thermodynamic properties of the black hole by analyzing its Hawking temperature. The Hawking temperature is defined as
$T=\frac{\kappa}{2\pi}=\frac{f^{\prime}\left(r_{H}\right)}{4\pi},$ (17)
where $\kappa$ represents the surface gravity at the event horizon. To evaluate this temperature, we first determine the black hole mass $m$ in terms of the horizon radius $r_{H}$ by solving the condition $f(r_{H})=0$. This leads to the following expression:
$m=\frac{r_{H}}{2}\left(1-8\pi\xi\eta^{2}\right)\left(1+\frac{Q^{3}}{r_{H}^{3}}\right).$ (18)
Substituting Eq. (18) into Eq. (17), we obtain the Hawking temperature:
$T_{H}=\frac{\left(1-8\pi\eta^{2}\xi\right)}{4\pi r_{H}}\frac{\left(r_{H}^{3}-2Q^{3}\right)}{\left(r_{H}^{3}+Q^{3}\right)}.$ (19)
For the Hawking temperature to remain both real and positive, two conditions must be satisfied. The first constraint stems from the nonlinear electromagnetic field, establishing a lower bound on the event horizon radius $r_{H}$ in terms of the magnetic charge $Q$:
$r_{H}\geq 2^{1/3}Q.$ (20)
The second constraint ensures the positivity of temperature when $\xi=1$, leading to an upper bound on the symmetry-breaking energy scale:
$\eta<\frac{1}{\sqrt{8\pi}}.$ (21)
These constraints highlight the interplay between the black hole horizon, charge, and the presence of the phantom global monopole, significantly influencing the thermodynamic behavior of the system.
A significant observation is that the Hawking temperature vanishes at $r_{H}=2^{1/3}Q$, indicating that the black hole ceases to emit Hawking radiation. In Figure 2, we depict the Hawking temperature of the nonlinear magnetic-charged black hole as a function of the horizon radius $r_{H}$ for three different cases: $\xi=-1,0,+1$.
(a) $\eta=0.05$
(b) $\eta=0.1$
(c) $\eta=0.15$
Figure 2: Hawking temperature $T_{H}$ as a function of the horizon radius $r_{H}$ for different values of the symmetry-breaking scale $\eta$. The black hole parameters are set to $Q=0.5$ and $m=1$. The effect of the phantom global monopole ($\xi=-1$) leads to an increase in temperature, whereas the ordinary global monopole ($\xi=+1$) lowers it.
From this plot, we observe that the Hawking temperature decreases as the horizon radius increases, meaning that larger black holes tend to have lower temperatures. When $\xi=-1$, the Hawking temperature is higher compared to the other cases for a given $r_{H}$, implying that the presence of the phantom global monopole amplifies the black hole’s temperature. Conversely, for $\xi=+1$, the temperature is the lowest among the three cases, confirming that while the phantom global monopole increases the black hole’s temperature, the ordinary global monopole reduces it. Furthermore, the figure illustrates that during the black hole’s evaporation process, the Hawking temperature initially rises, reaching a peak value $T_{\max}$ at a critical horizon radius $r_{H}^{\mathrm{c}}$. Beyond this point, the temperature rapidly declines, ultimately reaching zero. At this stage, the black hole stops radiating, and the pair-production processes occurring near $r_{H}^{\mathrm{c}}$ come to an end. Additionally, we note that as the energy scale of symmetry breaking $\eta$ increases, the Hawking temperature exhibits different behaviors in the two cases: it rises in the presence of the phantom global monopole but decreases when an ordinary global monopole is present.
We now turn our attention to the second thermodynamic quantity, entropy, which is defined by the relation:
$dS=\frac{dm}{T}.$ (22)
Using the expression for the black hole mass (18) and the Hawking temperature (19), we obtain the following expression for the entropy:
$S=\frac{A}{4}-\frac{4\pi^{3/2}Q^{3}}{\sqrt{A}},$ (23)
where $A=4\pi r_{H}^{2}$ represents the area of the event horizon. It is noteworthy that the entropy of the black hole remains the same for both the ordinary and phantom cases, suggesting that the presence of the phantom global monopole does not affect the black hole’s entropy.
Next, we investigate the thermodynamic stability of the black hole by analyzing its heat capacity. The heat capacity is defined as:
$C=T\left(\frac{\partial S}{\partial T}\right),$ (24)
and can be expressed, using Eqs. (19), (23), and (24), in the following form:
$C=-\frac{2\pi\left(r_{H}^{3}-2Q^{3}\right)\left(r_{H}^{3}+Q^{3}\right)^{2}}{r_{H}^{7}-10Q^{3}r_{H}^{4}-2Q^{6}r_{H}}.$ (25)
In Figure 3, we plot the heat capacity of the nonlinear magnetic-charged black hole as a function of the horizon radius $r_{H}$.
Figure 3: Heat capacity $C$ as a function of the horizon radius $r_{H}$ for the nonlinear magnetic-charged black hole. The discontinuity in $C$ indicates the presence of a second-order phase transition, separating an unstable region ($C<0$) from a stable region ($C>0$).
From this plot, we observe the presence of a discontinuity for each value of the magnetic charge parameter $Q$. Physically, this discontinuity signifies that the black hole undergoes a second-order phase transition. This transition separates an unstable phase, where $C<0$, from a stable phase, where $C>0$. Specifically, for smaller black holes (low $r_{H}$), the negative heat capacity indicates that they lose mass and evaporate due to Hawking radiation. However, beyond the critical point where $C$ changes sign, the black hole enters a stable phase, implying that it can remain in equilibrium with its surroundings.
To gain deeper insight into the thermal phase transition, we now analyze the Gibbs free energy $G$, which is defined as:
$G=m-TS.$ (26)
Substituting the expressions for the black hole mass (18), Hawking temperature (19), and entropy (23), we obtain the Gibbs free energy in the following form:
$G=\frac{\left(1-8\pi\eta^{2}\xi\right)\left(r_{H}^{6}+8Q^{3}r_{H}^{3}-2Q^{6}\right)}{4_{H}r_{H}^{2}\left(r_{H}^{3}+Q^{3}\right)}.$ (27)
It is important to emphasize that the Gibbs free energy plays a crucial role in determining the thermodynamic stability of the black hole. Specifically, when $G$ vanishes, the system reaches a critical threshold that separates stable and unstable phases. From Eq. (27), we find that this critical point occurs at:
$r_{H}=0.623\times Q.$ (28)
At this horizon radius, the Gibbs free energy tends to zero, indicating a potential phase transition in the black hole system.
In Figure 4, we plot the Gibbs free energy as a function of the event horizon radius $r_{H}$ for different values of the symmetry-breaking parameter $\eta$.
(a) $\eta=0.05$
(b) $\eta=0.1$
(c) $\eta=0.15$
Figure 4: Gibbs free energy $G$ as a function of the event horizon radius $r_{H}$ for different values of the symmetry-breaking parameter $\eta$. The plots illustrate the existence of a second-order phase transition, where $G$ changes sign at a critical radius $r_{H}^{c}$. For small black holes, $G>0$ suggests instability, whereas for larger black holes, $G<0$ indicates a stable phase.
The plots reveal the presence of a second-order phase transition. Notably, for small $r_{H}$, the Gibbs free energy is positive, suggesting an unstable phase, whereas for larger values of $r_{H}$, the free energy becomes negative, indicating a thermodynamically stable phase. The presence of the phantom global monopole ($\xi=-1$) shifts the critical point, influencing the black hole’s stability conditions.
4 Timelike Geodesics
In GR, test particles move along geodesics, which represent the natural trajectories dictated by the underlying spacetime curvature. When no external forces act upon a particle, its motion is governed solely by the spacetime geometry [65, 66, 67, 68]. The nature of a geodesic depends on the mass of the particle: massive particles traverse timelike geodesics, whereas massless particles, such as photons, follow null geodesics. The geodesic equation, which describes these trajectories, is given by
$\frac{d^{2}x^{\nu}}{d\lambda^{2}}+\Gamma_{\mu\sigma}^{\nu}\frac{dx^{\mu}}{d\lambda}\frac{dx^{\sigma}}{d\lambda}=0,$ (29)
where $\lambda$ represents an affine parameter along the geodesic, and $\Gamma_{\mu\sigma}^{\nu}$ are the Christoffel symbols associated with the spacetime metric.
For a spherically symmetric black hole spacetime, the geodesic motion simplifies, leading to the following system of differential equations:
$\ddot{t}+\frac{f^{\prime}(r)}{f(r)}\dot{r}\dot{t}=0,$ (30)
$\ddot{r}+\frac{f(r)}{2}\left(f^{\prime}(r)\dot{t}^{2}+\frac{\dot{r}}{f^{\prime}(r)}-2r\dot{\theta}^{2}-2r\sin^{2}\theta\dot{\phi}^{2}\right)=0,$ (31)
$\ddot{\theta}+2\frac{\dot{r}}{r}\dot{\theta}-\cos\theta\sin\theta\dot{\phi}^{2}=0,$ (32)
$\ddot{\phi}+2\frac{\dot{r}}{r}\dot{\phi}+2\cot\theta\dot{\theta}\dot{\phi}=0.$ (33)
[…]
$\ddot{t}+\frac{f^{\prime}(r)}{f(r)}\dot{r}\dot{t}=0,$ (34)
$\ddot{r}+\frac{f(r)}{2}\left(f^{\prime}(r)\dot{t}^{2}+\frac{\dot{r}}{f^{\prime}(r)}-2r\dot{\phi}^{2}\right)=0,$ (35)
$\ddot{\phi}+2\frac{\dot{r}}{r}\dot{\phi}=0.$ (36)
Geodesic motion obeys a conserved quantity that arises due to metric compatibility:
$g_{\mu\nu}\frac{dx^{\mu}}{d\lambda}\frac{dx^{\nu}}{d\lambda}=-1,$ (37)
Due to the presence of Killing vector fields associated with the metric symmetries, the geodesic equations admit two conserved quantities: the specific energy $E$ and the specific angular momentum $L$. These conserved quantities simplify the equations of motion:
$\dot{t}=\frac{E}{f(r)},$ (38)
$\dot{\phi}=\frac{L}{r^{2}}.$ (39)
By expressing the radial motion in terms of these conserved quantities, we obtain a first-order differential equation:
$\dot{r}^{2}+V_{\mathrm{eff}}(r)=E^{2},$ (40)
where the effective potential governing the radial motion is given by:
$V_{\mathrm{eff}}(r)=\left(1-8\pi\xi\eta^{2}-\frac{2mr^{2}}{r^{3}+Q^{3}}\right)\left(1+\frac{L^{2}}{r^{2}}\right).$ (41)
Eq. (40) describes a system analogous to the motion of a classical particle with energy $E^{2}$ in an effective one-dimensional potential $V_{\mathrm{eff}}(r)$. The nature of this potential determines key features of the geodesic motion, such as stable circular orbits, escape conditions, and critical impact parameters for light bending. The influence of the phantom field, nonlinear electrodynamics, and the global monopole manifests in modifications to this potential, altering the behavior of both massive and massless test particles.
On the other hand, For the motion of massive test particles, the effective potential plays a crucial role in determining orbital dynamics. It governs the radial motion and establishes the conditions for stable orbits. The effective potential for a test particle in this spacetime is given by:
$V_{\mathrm{eff}}(r)=\left(1-\frac{2mr^{2}}{r^{3}+Q^{3}}\right)+\frac{L^{2}}{r^{2}}-\frac{2m}{r^{3}+Q^{3}}L^{2}-8\pi\xi\eta^{2}\left(1+\frac{L^{2}}{r^{2}}\right).$ (42)
The terms in this expression contribute distinctively to the potential landscape: the second term represents a centrifugal barrier, the third term introduces relativistic corrections due to spacetime curvature, and the fourth term accounts for modifications arising from the presence of the phantom global monopole. Figure 5 illustrates the behavior of $V_{\mathrm{eff}}(r)$ for varying values of $\eta$.
(a) $\eta=0.05$
(b) $\eta=0.07$
(c) $\eta=0.1$
Figure 5: Variation of the effective potential $V_{\mathrm{eff}}(r)$ with different values of $\eta$, considering $Q=0.5$ and $m=1$.
Analyzing Eq. (42) asymptotically, we observe that at large distances, the effective potential approaches a limiting value:
$V_{\mathrm{eff}}(r)\rightarrow\left(1-8\pi\xi\eta^{2}\right)\quad\text{as}\quad r\to\infty.$ (43)
For a particle to escape to infinity, its total energy must satisfy the condition:
$E^{2}\geq 1-8\pi\xi\eta^{2}.$ (44)
Orbits that satisfy this condition are classified as unbound. In contrast, bound orbits require:
$E^{2}\leq 1-8\pi\xi\eta^{2}.$ (45)
Here, the term $1-8\pi\xi\eta^{2}$ represents the effective rest energy of the test particle in this modified spacetime.
To explore the nature of these orbits further, one can categorize the particle’s motion based on its energy relative to a critical value $E_{c}$:
  • $E>E_{c}$: The particle has enough energy to escape or fall directly into the singularity.
  • $E=E_{c}$: The particle undergoes unstable circular motion.
  • $E<E_{c}$: The particle follows a bound orbit, approaching a minimum radius before returning outward.
These behaviors are depicted in Figure 6, illustrating the variation of the effective potential under different values of the parameter $\xi$.
(a) $\xi=1$
(b) $\xi=0$
(c) $\xi=-1$
Figure 6: Effective potential for different values of $\xi$, with fixed $Q=0.5$, $\eta=0.05$, $L=10$, and $m=1$.
Furthermore, Figure 7 presents the trajectories of test particles in unstable circular orbits for varying angular momentum values.
(a) $L=10$
(b) $L=20$
(c) $L=30$
Figure 7: Unstable circular orbits for different angular momentum values, considering $Q=0.5$, $\eta=0.05$, and $m=1$.
For a qualitative analysis of circular orbits, it is necessary to examine the conditions under which the radial coordinate remains fixed. These conditions are derived by setting:
$\dot{r}=0\quad\Longrightarrow\quad V_{\mathrm{eff}}(r)=E^{2},$ (46)
$\ddot{r}=0\quad\Longrightarrow\quad V_{\mathrm{eff}}^{\prime}(r)=0,$ (47)
where $V_{\mathrm{eff}}^{\prime}(r)=\frac{\partial V_{\mathrm{eff}}(r)}{\partial r}$. Applying these conditions allows us to determine the specific energy $E$ and angular momentum $L$ required for stable circular orbits:
$L^{2}=\frac{r^{7}-2Q^{3}r^{4}}{Q^{6}-8\pi\eta^{2}\xi\left(Q^{3}+r^{3}\right)^{2}+2Q^{3}r^{3}+r^{6}-3r^{5}},$ (48)
$E^{2}=\frac{4\left(r^{4}-2Q^{3}r\right)^{2}}{\left(Q^{3}+r^{3}\right)^{2}\left(Q^{6}+2Q^{3}r^{3}+(r-3)r^{5}\right)-8\pi\eta^{2}\xi\left(Q^{3}+r^{3}\right)^{4}}.$ (49)
The stability of these orbits is determined by the second derivative of the effective potential:
[…]
The innermost stable circular orbit (ISCO) is found by imposing the additional condition:
$\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}=0.$ (52)
A deeper understanding of orbital motion can be gained by analyzing circular orbits, where the radial coordinate $r$ remains constant. To sustain a circular trajectory, a test particle must satisfy two conditions. The first ensures that the radial velocity vanishes:
$\dot{r}=0\quad\Longrightarrow\quad V_{\mathrm{eff}}(r)=E^{2}.$ (53)
The second condition ensures that there is no radial acceleration:
$\ddot{r}=0\quad\Longrightarrow\quad V_{\mathrm{eff}}^{\prime}(r)=0,$ (54)
where the prime denotes differentiation with respect to $r$. Using these conditions, one can determine the specific angular momentum $L$ and specific energy $E$ required for a test particle to maintain a circular orbit:
$L^{2}=\frac{r^{7}-2Q^{3}r^{4}}{Q^{6}-8\pi\eta^{2}\xi(Q^{3}+r^{3})^{2}+2Q^{3}r^{3}+r^{6}-3r^{5}},$ (55)
$E^{2}=\frac{4(r^{4}-2Q^{3}r)^{2}}{(Q^{3}+r^{3})^{2}(Q^{6}+2Q^{3}r^{3}+(r-3)r^{5})-8\pi\eta^{2}\xi(Q^{3}+r^{3})^{4}}.$ (56)
The stability of these circular orbits is determined by the second derivative of the effective potential:
[…]
The innermost stable circular orbit (ISCO) is of particular significance as it marks the transition between stable and unstable motion. At the ISCO, the maximum and minimum of the effective potential merge, imposing the additional condition:
$\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}=0.$ (59)
By substituting Eq. (42) into this condition, we derive the ISCO constraint:
$\left.rf^{\prime\prime}(r)f(r)+3f(r)f^{\prime}(r)-2rf^{\prime}(r)^{2}\right|_{r=r_{\mathrm{ISCO}}}=0.$ (60)
Expanding this equation explicitly, we obtain:
$0\left.11Q^{3}r^{4}(1-8\pi\eta^{2}\xi)-8Q^{6}r(1-8\pi\eta^{2}\xi)+r^{6}(r-6-8\pi\eta^{2}\xi r)\right|_{r=r_{\mathrm{ISCO}}}=$ 0 . (61)
To gain further insight into the ISCO properties, we numerically evaluate the ISCO radius $r_{\mathrm{ISCO}}$, as well as the corresponding angular momentum $L_{\mathrm{ISCO}}$ and energy $E_{\mathrm{ISCO}}$, for different values of $\eta$ and $\xi$. The results are summarized in Table 1.
$\xi=1$ $\xi=-1$
$\eta$ $r_{\mathrm{ISCO}}$ $L_{\mathrm{ISCO}}$ $E_{\mathrm{ISCO}}$ $r_{\mathrm{ISCO}}$ $L_{\mathrm{ISCO}}$ $E_{\mathrm{ISCO}}$
$0.01$ 5.97664 3.46678 0.0793058 5.94608 3.44931 0.0799223
$0.05$ 6.36837 3.69102 0.0720584 5.6015 3.25243 0.0874702
$0.09$ 7.50929 4.34573 0.0562133 4.92859 2.8693 0.106192
$0.12$ 9.38748 5.42642 0.0401827 4.3324 2.5321 0.129227
Table 1: Numerical values of the ISCO parameters $r_{\mathrm{ISCO}}$, $L_{\mathrm{ISCO}}$, and $E_{\mathrm{ISCO}}$ for test particles under different values of $\eta$ and $\xi$.
Another key aspect of orbital stability that we analyze is the principal Lyapunov exponent, which characterizes the divergence or convergence of neighboring trajectories in phase space. This exponent quantifies how small perturbations affect the stability of circular orbits. A positive Lyapunov exponent indicates an unstable orbit where nearby trajectories diverge over time, whereas a negative Lyapunov exponent corresponds to a stable orbit with converging trajectories. The Lyapunov exponent $\lambda_{L}$ is defined as:
$\lambda_{L}=\sqrt{-\frac{1}{2\dot{t}^{2}}\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}}.$ (62)
By substituting Eqs. (38) and (42) into the above expression, we derive the explicit form of $\lambda_{L}$ for a nonlinear magnetic-charged black hole with a phantom global monopole:
$\lambda_{L}$ $=$ $\frac{1}{E}\left(1-8\pi\eta^{2}\xi-\frac{2mr^{2}}{Q^{3}+r^{3}}\right)\Biggr{[}\frac{2L^{2}}{r^{3}}\left(\frac{6mr^{4}}{(Q^{3}+r^{3})^{2}}-\frac{4mr}{Q^{3}+r^{3}}\right)-\frac{3L^{2}}{r^{4}}\left(1-8\pi\eta^{2}\xi-\frac{2mr^{2}}{Q^{3}+r^{3}}\right)$ (63)
$+\frac{m}{Q^{3}+r^{3}}\left(1+\frac{L^{2}}{r^{2}}\right)\left(1-\frac{9r^{3}}{(Q^{3}+r^{3})}+\frac{9r^{6}}{(Q^{3}+r^{3})^{2}}\right)\Biggr{]}^{1/2}.$
In Figure 8, we illustrate the radial dependence of the Lyapunov exponent for test particles orbiting a nonlinear magnetic-charged black hole with a phantom global monopole. The behavior of $\lambda_{L}$ provides crucial insight into the stability of circular orbits.
(a) $\eta=0.12$
(b) $\eta=0.1$
(c) $\eta=0.08$
Figure 8: Radial dependence of the Lyapunov exponent $\lambda_{L}$ for test particle motion around a nonlinear magnetic-charged black hole with a phantom global monopole. The parameters used are $Q=0.5$, $L=5$, $E=1$, and $m=1$.
From Figure 8, we observe that as the energy scale of symmetry breaking, represented by $\eta$, decreases, the region allowing stable circular orbits expands. This suggests that a weaker global monopole field enhances orbital stability by reducing the disruptive effects on geodesic motion. Additionally, the Lyapunov exponent curves intersect the horizontal axis at specific radii, marking the location of the ISCO. These bifurcation points indicate a transition between stable and unstable circular orbits. Furthermore, as $\eta$ decreases, the ISCO radius shifts outward, implying that particles in lower-energy monopole fields require a larger orbital radius to maintain stability. This trend highlights the significant role of the global monopole in shaping the phase-space dynamics of test particle motion in modified gravity scenarios.
5 Quasinormal Modes
Quasinormal modes characterize the response of a black hole to external perturbations and provide crucial insights into its stability and observational signatures in gravitational wave astronomy [69, 70]. The spectrum of these modes depends on the black hole parameters and surrounding matter fields, making them valuable probes of deviations from general relativity.
In the presence of a global monopole, the quasinormal spectrum is modified due to the conical deficit introduced by the monopole field. Prior studies have examined quasinormal ringing in such backgrounds for scalar [71, 72, 73], Dirac [74], Maxwell [75], and spinor fields [76]. These works show that the presence of a monopole shifts the real and imaginary parts of the QNM frequencies, affecting the decay rate of perturbations. More recent research in modified gravity scenarios, such as Einstein–bumblebee gravity, has further explored the influence of global monopoles on black hole stability and QNM spectra [77, 78, 79].
Additionally, NED introduces further modifications to the QNM spectrum by altering the effective potential governing perturbations. These effects have been investigated in contexts such as Einstein-Gauss-Bonnet gravity [80], AdS black holes with Coulomb-like corrections [81], and charged black holes with nonlinear electrodynamics-induced modifications [82, 83]. Studies have also examined how these effects influence black hole shadows and grey-body factors [84, 85], while additional work on Euler-Heisenberg black holes suggests notable shifts in the QNM spectrum [86].
In this section, we analyze the quasinormal modes of a nonlinear magnetic-charged black hole with a phantom global monopole. Using the sixth-order WKB approximation, we compute the mode spectrum and compare our results with the Pöschl-Teller potential approach to assess accuracy. Our goal is to determine how the interplay between the phantom global monopole and nonlinear electrodynamics influences the black hole’s dynamical response and stability under scalar perturbations.
Let us consider the dynamics of a massless scalar field propagating in the black hole background, governed by the wave equation:
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\psi\right)=0.$ (64)
Applying the method of separation of variables, we express the field as:
$\psi_{n,\ell,\mu}(r,\theta,\phi,t)=\frac{R_{n,\ell}(r)}{r}\mathcal{Y}_{\ell,\mu}(\theta,\phi)e^{-i\omega t},$ (65)
where $\mathcal{Y}_{\ell,\mu}(\theta,\phi)$ are the spherical harmonics describing the angular dependence, and $R_{n,\ell}(r)$ is the radial function satisfying the equation:
$f(r)\frac{d}{dr}\left[f(r)\frac{d}{dr}R_{n,\ell}(r)\right]+\Big{(}\omega^{2}-\mathcal{V}_{\mathrm{eff}}(r)\Big{)}R_{n,\ell}(r)=0.$ (66)
Here, the effective potential governing the scalar perturbations is given by:
$\mathcal{V}_{\mathrm{eff}}(r)=f(r)\left(\frac{\ell(\ell+1)}{r^{2}}+\frac{1}{r}\frac{d}{dr}f(r)\right).$ (67)
To further simplify the radial equation, we introduce the tortoise coordinate $x$:
$x=\int\frac{dr}{f(r)},$ (68)
which transforms the radial equation into a Schrödinger-like form:
$\frac{d^{2}}{dx^{2}}R_{n,\ell}(x)+\Big{(}\omega^{2}-\mathcal{V}_{\mathrm{eff}}(x)\Big{)}R_{n,\ell}(x)=0.$ (69)
The behavior of the quasinormal modes is determined by imposing physically motivated boundary conditions on Eq. (69). These conditions ensure purely ingoing waves at the black hole horizon and purely outgoing waves at spatial infinity, leading to:
$R_{n,\ell}(x)\thickapprox e^{\pm i\omega x}\quad\text{as}\quad x\longrightarrow\pm\infty.$ (70)
The effective potential $\mathcal{V}_{\mathrm{eff}}(x)$ determines the QNM spectrum, as it represents the potential barrier through which perturbations must tunnel before being absorbed by the black hole. Its shape and peak height influence the stability and damping rates of the modes. In the following, we analyze the numerical methods used to compute the QNM spectrum and discuss the effects of the phantom global monopole and nonlinear electrodynamics on the stability of the black hole.
5.1 Computation of QNMs using the WKB Approximation
To compute the quasinormal frequencies, we employ the WKB approximation, a widely used semi-analytical method in black hole physics [89, 90, 87, 88, 94, 95, 96, 86, 97, 98, 93, 91, 92]. Originally introduced at first order in [99], the method was later extended to higher orders, including second and third order in [100] and up to sixth order in [101]. While higher-order corrections enhance precision, their effectiveness depends on the potential shape and parameter regime, as discussed in the reviews [70, 102].
In this work, we use the sixth-order WKB formula from [101], given by:
$\frac{i\left(\omega_{n}^{2}-V_{0}\right)}{\sqrt{-2V_{0}^{\prime\prime}}}-\sum_{j=2}^{6}\Phi_{j}=n+\frac{1}{2},$ (71)
where $V_{0}$ represents the peak of the effective potential at $x_{0}$, the notation (′′) denotes the second derivative with respect to the tortoise coordinate, $\Phi_{j}$ accounts for higher-order corrections, and $n$ is the overtone number. The resulting quasinormal frequencies $\omega_{n}$ form a discrete spectrum, where the real part corresponds to the oscillation frequency, and the imaginary part determines the damping rate.
Since Eq. (71) involves nontrivial dependencies on physical parameters, we numerically compute the QNM frequencies for different parameter values. However, as the WKB approximation is most reliable when $\ell>n$, we restrict our analysis to scalar field perturbations satisfying this condition, corresponding to the low-lying quasinormal modes. Selected results for these modes are presented in Table 2
$\ell$ $n$ $\omega_{WKB}(\eta=0.12)$ $\omega_{WKB}(\eta=0.09)$ $\omega_{WKB}(\eta=0.06)$ $\omega_{WKB}(\eta=0.03)$
1 0 0.145742-i 0.0394710 0.205796-i 0.0614642 0.253484-i 0.0800628 0.283985-i 0.0923490
2 0 0.244576-i 0.0392214 0.342822-i 0.0610029 0.420085-i 0.0794028 0.469184-i 0.0915496
1 0.237964-i 0.119079 0.331583-i 0.185690 0.404717-i 0.242115 0.451016-i 0.279432
3 0 0.343058-i 0.0391535 0.479908-i 0.0608754 0.587231-i 0.0792188 0.655308-i 0.0913261
1 0.338235-i 0.118187 0.471645-i 0.184011 0.575871-i 0.239683 0.641833-i 0.276461
2 0.329068-i 0.199332 0.456143-i 0.311136 0.554751-i 0.405941 0.616916-i 0.468663
4 0 0.441421-i 0.0391258 0.617006-i 0.0608231 0.754549-i 0.0791432 0.841728-i 0.0912343
1 0.437638-i 0.117818 0.610504-i 0.183311 0.745590-i 0.238664 0.831087-i 0.275217
2 0.430296-i 0.197808 0.597982-i 0.308271 0.728427-i 0.401792 0.810766-i 0.463613
3 0.419844-i 0.279875 0.580390-i 0.437128 0.704537-i 0.570554 0.782633-i 0.658860
Table 2: QNMs of the nonlinear magnetic-charge BH with ordinary global monopole ($\xi=1$), calculated using the 6th-order WKB approximation method, for $m=1$ and $Q=0.5$.
and in Table 3.
$\ell$ $n$ $\omega_{WKB}(\eta=0.12)$ $\omega_{WKB}(\eta=0.09)$ $\omega_{WKB}(\eta=0.06)$ $\omega_{WKB}(\eta=0.03)$
1 0 0.483648-i 0.176601 0.395847-i 0.139083 0.337947-i 0.114637 0.305100-i 0.100997
2 0 0.785554-i 0.174813 0.647424-i 0.137734 0.555511-i 0.113578 0.503043-i 0.100098
1 0.747231-i 0.535822 0.618247-i 0.421576 0.532151-i 0.347187 0.482875-i 0.305711
3 0 1.09191-i 0.174313 0.901696-i 0.137354 0.774781-i 0.113280 0.702201-i 0.0998456
1 1.06303-i 0.528917 0.879837-i 0.416441 0.757366-i 0.343204 0.687212-i 0.302355
2 1.01100-i 0.900111 0.840080-i 0.707800 0.725436-i 0.582622 0.659590-i 0.512855
4 0 1.39975-i 0.174109 1.15685-i 0.137198 0.994608-i 0.113158 0.901750-i 0.0997421
1 1.37678-i 0.525991 1.13952-i 0.414273 0.980824-i 0.341528 0.889903-i 0.300945
2 1.33358-i 0.888399 1.10672-i 0.699087 0.954632-i 0.575858 0.867323-i 0.507150
3 1.27527-i 1.26657 1.06207-i 0.995671 0.918684-i 0.819346 0.836173-i 0.721089
Table 3: QNMs of the nonlinear magnetic-charge BH with phantom global monopole ($\xi=-1$), calculated using the 6th-order WKB approximation method, for $m=1$ and $Q=0.5$.
The numerical results given in Tables 2 and 3 underlines the several key results:
  • The imaginary part of the frequencies is consistently negative, confirming the stability of scalar field perturbations in this black hole background. This stability arises from the positive potential barrier outside the event horizon, which prevents perturbations from growing unbounded.
  • As $\ell$ increases for a fixed $\eta$, the real part of the frequency also increases. This trend is expected, as larger angular momentum states correspond to higher oscillation frequencies.
  • For a given overtone number $n$ and angular momentum $\ell$, increasing the symmetry-breaking parameter $\eta$ generally increases the real part of the frequency in the case of a black hole with an ordinary global monopole. However, for a black hole with a phantom global monopole, the opposite trend is observed—the real part of the QNMs decreases with increasing $\eta$. This distinction suggests that the phantom monopole modifies the effective gravitational potential in a way that reduces oscillation frequencies.
Figure 9 illustrates an example of QNMs computed by using the sixth-oreder WKB formula for scalar perturbations with $\eta=0.12$, $\xi=-1$ and $Q=0.3$.
(a) $\ell=4$, $n=2$
(b) $\ell=4$, $n=6$
Figure 9: The variation of QNMs for scalar perturbations in a nonlinear magnetic-charged black hole with a phantom global monopole for $\eta=0.12$, $\xi=-1$ and $Q=0.3$. The plot illustrates the dependence of QNMs on the overtone number $n$ and angular momentum $\ell$, highlighting the increasing deviation of the WKB approximation for higher overtones.
As evident from the figure, the accuracy of the WKB formula decreases for scalar perturbations when $n\geq l$, whereas the results converge more reliably for cases where $l>n$.
Figure 10 illustrates the changes in the QNMs as a function of $n$ for both ordinary and phantom global monopoles of a nonlinear magnetic-charged black hole.
(a) $\xi=1$
(b) $\xi=-1$
(c) $\xi=1$
(d) $\xi=-1$
Figure 10: QNM frequencies as a function of overtone number $n$ for $\ell=10$, $m=1$, $\eta=0.12$. Higher overtones exhibit lower real frequencies and stronger damping. The effect of the global monopole parameter $\xi$ is also visible.
Figure 10 illustrates the dependence of quasinormal mode frequencies on the overtone number $n$ for scalar perturbations. As $n$ increases, the real part of the frequency decreases, leading to lower oscillation frequencies, while the imaginary part increases, indicating stronger damping. This implies that higher overtones decay more rapidly. Additionally, the presence of a global monopole significantly affects the QNM spectrum, with notable differences between the ordinary ($\xi=1$) and phantom ($\xi=-1$) cases. The results highlight how the monopole field influences the stability and decay of perturbations in the black hole spacetime.
5.2 Pöschl-Teller potential approximation
This approach consists of approximating the potential $\mathcal{V}_{\mathrm{eff}}\left(x\right)$ with a similarly shaped potential, for which the solutions of Eq. (69) can be computed analytically. The Pöschl-Teller (PT) potential satisfies this condition and has the following form:
$V_{\mathrm{PT}}\left(x\right)=\frac{V_{0}}{\cosh^{2}\left(\frac{x}{b}\right)},$ (72)
where
$b=\frac{1}{\sqrt{-\frac{1}{2V_{0}}\frac{d^{2}V\left(x_{0}\right)}{dx^{2}}}}.$ (73)
We now adopt the approach outlined in [103] to illustrate how the QNM frequencies can be computed analytically for this approximate potential. The key strategy is to transform equation (69) into a form that resembles the Schrödinger wave equation, such that the solutions of (69) that satisfy the boundary condition (70) correspond to bound states of the transformed equation. As is well known, the bound states of the Schrödinger equation can be calculated analytically. After computing them, the inverse transformation is applied to obtain the QNMs. The QNMs for the potential $V_{\mathrm{PT}}\left(x\right)$ can be calculated analytically and are expressed as follows:
$\omega_{\mathrm{PT}}=\frac{1}{b}\left[-i\left(n+\frac{1}{2}\right)\pm\sqrt{V_{0}b^{2}-\frac{1}{4}}\right].$ (74)
Now, we compute the quasinormal frequencies for scalar field perturbations for different values of $\eta$, utilizing Eq. (74). The results are tabulated in Table 4.
$\eta$ $\omega_{WKB}(\xi=1)$ $\omega_{PT}(\xi=1)$ $\omega_{WKB}(\xi=-1)$ $\omega_{PT}(\xi=-1)$
0.01 0.877073 -i 0.0956334 0.297450-i 0.100073 0.883812 - i 0.0965986 0.299827 -i 0.101102
0.02 0.866999 -i 0.0941947 0.293899-i 0.0985414 0.893955 -i 0.0980555 0.303406 -i 0.102654
0.03 0.850299 -i 0.0918209 0.288018-i 0.0960147 0.910951 -i 0.100508 0.309409 -i 0.105268
0.04 0.827113 -i 0.0885483 0.279864-i 0.0925337 0.934935 -i 0.103992 0.317891 -i 0.108984
0.05 0.797636 -i 0.0844276 0.269515-i 0.0881547 0.966094 -i 0.108558 0.328929 -i 0.113859
0.06 0.762124 -i 0.0795242 0.257076-i 0.0829500 1.00467 -i 0.114271 0.342624 -i 0.119965
0.07 0.720893 -i 0.0739178 0.242674-i 0.0770077 1.05094 -i 0.121211 0.359095 -i 0.127392
0.08 0.674327 -i 0.0677029 0.226461-i 0.0704316 1.10525 -i 0.129471 0.378483 -i 0.136244
0.09 0.622884 -i 0.0609885 0.208617-i 0.0633412 1.16797 -i 0.139161 0.400949 -i 0.146644
0.1 0.567097 -i 0.0538981 0.189349-i 0.0558712 1.23952 -i 0.150401 0.426674 -i 0.158731
Table 4: Comparison of QNMs of the scalar field by using PT approximation method, and 6th-order WKB approximation method for $\ell=4$, $n=0$, and $m=1$, $Q=0.2$.
Comparing the PT approximation and the WKB method, we observe that WKB consistently predicts higher real frequencies and stronger damping. Additionally, the presence of a phantom monopole results in larger $\text{Re}(\omega)$ and increased damping compared to the ordinary monopole case, highlighting the influence of $\xi$ on the stability and decay of perturbations.
6 Grey-body factors
In this section, we delve into the grey-body factors which are closely related QNMs [104, 105, 106, 107, 108, 109]. The grey-body factor represents the probability that radiation emitted close to the event horizon will escape to a distant observer rather than being deflected back by the black hole’s gravitational potential [110]. This potential is influenced by the spacetime geometry of the black hole. For the scattering problem, we consider the following wave equation
$\frac{d^{2}}{dx^{2}}R(x)+\big{(}\Omega-\mathcal{V}_{\mathrm{eff}}(x)\big{)}R(x)=0.$ (75)
The boundary conditions used to solve this wave equation are determined by the behavior of the field both close to the event horizon and at infinity.
  • Near the event horizon , $x\rightarrow-\infty$, the field is entirely ingoing
    $R(x)\simeq\mathbf{T}e^{-i\Omega x},$ (76)
    where $\mathbf{T}$ is the transmission coefficient.
  • At infinity, $x\rightarrow+\infty$, the field consists of a combination of both ingoing and outgoing waves
    $R(x)\simeq e^{-i\Omega x}+\mathbf{R}e^{i\Omega x},$ (77)
    where $\mathbf{R}$ is the reflection coefficient.
These boundary conditions are employed to calculate the intensity of Hawking radiation [111, 112, 113], or for studying various radiation-related phenomena near black holes, such as superradiance [114, 115]. In the context of black hole radiation, the transmission coefficient, $\mathbf{T}$, is also known as the grey-body factor. This factor depends on the black hole’s spacetime geometry and the properties of the radiation,
$\Gamma_{\ell}\left(\Omega\right)=\left|\mathbf{T}\right|^{2}=1-\left|\mathbf{R}\right|^{2}.$ (78)
Grey-body factors and quasinormal modes can be computed using the higher-order WKB. The formulas were established for spherically symmetric and asymptotically flat black holes by applying the well-known WKB method to calculate grey-body factors,
$\Gamma_{\ell}\left(\Omega\right)=\frac{1}{1+e^{i2\pi\mathcal{K}}}.$ (79)
Here $\mathcal{K}$ depends on the values of the effective potential and its derivatives up to the 2i-th order for the third-order WKB method [99, 100]. As noted in [105], the function $\mathcal{K}$ can also be determined through an alternative approach, leveraging its connection to the fundamental mode and the first overtone of the quasinormal spectrum. Therefore, we have:
$i\mathcal{K}$ $=$ $\frac{\Omega^{2}-\mathrm{Re}\left(\omega_{0}\right)^{2}}{4\mathrm{Re}\left(\omega_{0}\right)\mathrm{Im}\left(\omega_{0}\right)}\left(1+\frac{\big{(}\mathrm{Re}\left(\omega_{0}\right)-\mathrm{Re}\left(\omega_{1}\right)\big{)}^{2}}{32\mathrm{Im}\left(\omega_{0}\right)^{2}}-\frac{3\mathrm{Im}\left(\omega_{0}\right)-\mathrm{Im}\left(\omega_{1}\right)}{24\mathrm{Im}\left(\omega_{0}\right)}\right)$ (80)
-- $\frac{\Big{(}\mathrm{Re}\left(\omega_{0}\right)-\mathrm{Re}\left(\omega_{1}\right)\Big{)}}{16\mathrm{Im}\left(\omega_{0}\right)}-\frac{\left(\Omega^{2}-\mathrm{Re}\left(\omega_{0}\right)^{2}\right)^{2}}{16\mathrm{Re}\left(\omega_{0}\right)^{3}\mathrm{Im}\left(\omega_{0}\right)}\Bigg{(}1-\frac{\mathrm{Re}\left(\omega_{0}\right)\big{(}\mathrm{Re}\left(\omega_{0}\right)-\mathrm{Re}\left(\omega_{1}\right)\big{)}}{4\mathrm{Im}\left(\omega_{0}\right)^{2}}\Bigg{)}$
$+$ $\frac{\left(\Omega^{2}-\mathrm{Re}\left(\omega_{0}\right)^{2}\right)^{3}}{32\mathrm{Re}\left(\omega_{0}\right)^{5}\mathrm{Im}\left(\omega_{0}\right)}\Bigg{[}1+\frac{\mathrm{Re}\left(\omega_{0}\right)\left(\mathrm{Re}\left(\omega_{0}\right)-\mathrm{Re}\left(\omega_{1}\right)\right)}{4\mathrm{Im}\left(\omega_{0}\right)^{2}}$
$+$ $\mathrm{Re}\left(\omega_{0}\right)^{2}\left(\frac{\left(\mathrm{Re}\left(\omega_{0}\right)-\mathrm{Re}\left(\omega_{1}\right)\right)^{2}}{16\mathrm{Im}\left(\omega_{0}\right)^{4}}-\frac{3\mathrm{Im}\left(\omega_{0}\right)-\mathrm{Im}\left(\omega_{1}\right)}{12\mathrm{Im}\left(\omega_{0}\right)}\right)\Bigg{]}+\mathcal{O}\left(\frac{1}{\ell^{3}}\right).$
Here $\omega_{0}$ is the the fundamental mode, and $\omega_{1}$ is the first overton. Figure 11 illustrates the grey-body factors as a function of frequency $\Omega$ for a nonlinear magnetic-charged black hole with both ordinary ($\xi=1$) and phantom ($\xi=-1$) global monopoles.
(a) $\xi=1$
(b) $\xi=-1$
Figure 11: Grey-body factors for $\ell=3$ and $Q=0.1$
The grey-body factor represents the probability that radiation emitted near the event horizon escapes to infinity rather than being reflected back by the potential barrier. As seen in the plot, the grey-body factor increases with frequency, meaning that higher-energy modes have a higher probability of transmission. Additionally, the presence of a phantom global monopole ($\xi=-1$) leads to a lower transmission coefficient compared to the ordinary case, indicating that the phantom field modifies the effective potential in a way that enhances reflection and suppresses radiation escape.
In Figure 12, we examine the dependence of the grey-body factors on the symmetry-breaking parameter $\eta$ for black holes with both ordinary and phantom global monopoles.
(a) $\xi=1$
(b) $\xi=-1$
Figure 12: Grey-body factors for $\ell=3$ and $\eta=0.1$
The results indicate that as $\eta$ increases, the transmission probability decreases, leading to a suppression of radiation escaping to infinity. This effect is more pronounced in the phantom case ($\xi=-1$), suggesting that stronger symmetry breaking increases the effective potential barrier surrounding the black hole. This result aligns with our findings in thermodynamics, where increasing $\eta$ leads to modifications in black hole stability and temperature, further influencing the radiation emission process.
Figure 13 explores the effect of the nonlinear magnetic charge $Q$ on the grey-body factors.
(a) $\eta=0.05$
(b) $\eta=0.11$
Figure 13: Grey-body factors for $\ell=3$, $\eta=0.1$ and $Q=0.2$
As seen in the figure, increasing $Q$ leads to a reduction in the transmission coefficient, meaning that a stronger magnetic charge suppresses the escape of radiation. This behavior is expected since a larger $Q$ enhances the curvature effects and modifies the effective potential, making it more difficult for radiation to overcome the potential barrier. Notably, this effect is consistent across both ordinary and phantom monopole cases, reinforcing the role of nonlinear electrodynamics in shaping black hole radiation characteristics.
7 Conclusion
In this work, we investigated the properties of a nonlinear magnetic-charged black hole in the presence of a phantom global monopole. By incorporating nonlinear electrodynamics (NED) and exotic scalar fields, we derived an exact black hole solution and analyzed its thermodynamic and dynamical properties. Our findings reveal significant deviations from standard general relativity (GR), providing potential observational signatures for modified gravity theories.
We examined the black hole’s thermodynamic properties, including its Hawking temperature, entropy, and heat capacity. The presence of the phantom global monopole was shown to modify the thermodynamic stability, leading to novel phase structures. Notably, the Hawking temperature exhibits a dependence on the monopole parameter, influencing the evaporation process of the black hole. The heat capacity analysis indicated the presence of a second-order phase transition, suggesting regions of stability and instability in the black hole’s thermodynamic phase space.
We also analyzed the geodesic structure of test particles, exploring their motion in the modified black hole background. Furthermore, we studied the quasinormal modes (QNMs) of the black hole, employing both the sixth-order WKB approximation and the Pöschl-Teller potential method. The QNM spectrum provided insights into the stability of the system under perturbations. Our results showed that the presence of nonlinear electrodynamics and the phantom global monopole modifies the real and imaginary parts of the QNM frequencies, affecting the damping times of perturbations.
A key addition to our study was the analysis of grey-body factors, which describe the probability of radiation escaping the black hole’s gravitational potential. We found that the presence of the phantom global monopole reduces the transmission coefficient, leading to a suppression of escaping radiation. This effect was further influenced by the symmetry-breaking parameter and the magnetic charge, reinforcing the role of nonlinear electrodynamics in shaping black hole radiation properties.
Overall, our findings highlight the intricate interplay between nonlinear electrodynamics, phantom fields, and global monopoles in black hole physics. The modifications introduced by these exotic fields have direct implications for astrophysical observations, including black hole shadows, gravitational waves, and Hawking radiation. Future work could extend this analysis by exploring rotating solutions, higher-dimensional generalizations, or the impact of additional modified gravity corrections.
Conflict of Interests
The authors declare no such conflict of interest.
Data Availability Statement
No data were generated or analyzed in this study.
Acknowledgments
This work is supported by the Ministry of Higher Education and Scientific Research, Algeria under the code: B00L02UN040120230003. B. C. L. is grateful to Excellence Project PřF UHK 2205/2025-2026 for the financial support.